OpenStudy (anonymous):

Sec^2x tan x=2tanx Solve over the interval 0,360 degrees

OpenStudy (anonymous):

what is sec in terms of tan?

OpenStudy (anonymous):

are u only seeking the answer? just ask if u dont get my guiding question...

OpenStudy (anonymous):

hello~~ here, do u know what secant is? what is it equal to in terms of tan?? this will help u simplify

OpenStudy (anonymous):

sec(x) = 1/[sin(pi/2 - x)] = 1/cos(x) = tan(x)/sin(x), sorry, i was busy doing my school requirements. @jolee